OpenStudy (anonymous):

Multiply. (x - 4)(x2 + 5x + 6) can someone confirm my answer?

OpenStudy (anonymous):

x3 + x2 - 14x - 24? thats what I got

OpenStudy (anonymous):

ur saying x to the power of 2 correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yah i got the same answer as itsjustme

OpenStudy (anonymous):

thanks

hero (hero):

(x - 4)(x2 + 5x + 6) = x^2(x -4) + 5x(x-4) + 6(x - 4) = x^3 - 4x^2 + 5x^2 - 20x + 6x - 24 = x^3 + x^2 -14x - 24